OpenStudy (anonymous):

if f(x)=-3x+3, then f^-1(x)=

OpenStudy (thefurball):

The inverse of something is just 1 divided by that thing. That applies to functions too. So, \[f(x)^{-1} = \frac{ 1 }{ f(x) }\] Can you figure it out now?

OpenStudy (anonymous):

no not actually

OpenStudy (thefurball):

Ok, I will walk you through it. It is just\[f(x)^{-1} = \frac{ 1 }{ -3x + 3 }\]

OpenStudy (anonymous):

so do i divid that? or no

OpenStudy (anonymous):

@Data_LG2 can you help me.

OpenStudy (anonymous):

To get the inverse of a function, 1. Let f(x) = y 2. Rewrite your function in terms of y like this y= 2x + 3 3. Switch the places of x and y, so it will be like this x= 2y + 3 4. Re-arrange the new equation terms of y, so it will be like this: x = 2y + 3 (subtract both sides by 3) x-3 = 2y (divide both sides by 2) (x-3)/2 = y 5. Now that will be your inverse function. Rewrite it in terms od f(x)^-1 it will be \(\sf \large f(x)^{-1}= \frac{(x-3)}{2}\) Now do it with your problem. I'll check it later (: